# Interactive Inference: A Neuromorphic Theory of Human-Computer Interaction

### v4

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXbooktitle: ACM Transactions on Computer-Human Interactionisbn: 978-1-4503-XXXX-X/18/06
@@ Introduction
Shannon went on to define that the information capacity $C$ of the transmission of a message through an analog channel has a maximum that is the (binary) logarithm of the ratio between the signal strength $S$ and the amount of Gaussian noise NN in the channel, multiplied by some constant $b$, as expressed by the Shannon-Hartley Theorem:
@@
Where $b$ is some empirically derived bandwidth parameter. This equation describes the capacity of information sent through the channel in terms of the signal-to-noise ratio. Note that without the negative sign, the SNR represents the reciprocal of the probability of a transmission. A high signal allows for more uncertainty to be carried, as does a lower noise. This means the very nature of information is statistical, with underlying equations based on (e.g., normal) distributions. This is important because humans, unlike traditional computer algorithms, operate in a probabilistic environment. That is, their exact behaviour cannot be computed, however, it can be modeled using (normal) probability distributions. Using this framework, Card, Moran and Newell (Card et al., 1983) developed the Model Human Processor (MHP), a comprehensive set of heuristics for predicting human performance in simple computer tasks. Statistical elements to these models were derived from empirical observations such as those performed by Hick (Hick, 1952) and Fitts (Fitts, 1954). They demonstrated that human performance in certain tasks indeed seems to follow a similar pattern to the Shannon-Hartley Theorem: a binary logarithm of the signal-to-noise ratio in the task. E.g., Hick’s Law describes the response time RTRT for selecting from a number of alternatives $n$, plus 1:
@@
Here, $b$ is some empirically derived constant that models the time per unit of difficulty, the log component of the task. Response time is derived by multiplying the amount of information in the task by this empirical constant. Fitts (Fitts, 1954) derived a very similar law describing the movement time MTMT of a hand towards a target along a single dimensional trajectory AA with a target width of $W$ (specific equation from (MacKenzie, 1992)):
@@
Instead of framing human processing as a sequential computer processor, McClelland and Rumelhart’s parallel processing model (McClelland and Rumelhart, 1981) described perception as an interaction of top-down (cognitive/semantic) and bottom-up (sensory) elements. Rao and Ballard (Rao and Ballard, 1999) further elucidated this approach by suggesting that the brain builds up statistical knowledge of its environment, i.e., incoming messages, over time, and that this model is used generatively, i.e., top-down, to predict the subsequent messages. The error between the prediction and the messages would lead to an update of the prediction. Prediction is convenient for more than one reason: not only does it allow humans to plan actions in their environment based on relevant information, it serves to compress the information content of their environment by processing only those messages that are surprising: i.e., not, or insufficiently predicted. Predictive coding theory explains how the human brain can work so efficiently and explains the mechanisms behind attention and the scheduling of cognitive resources based on the amount of surprise in their predictions (Baldi and Itti, 2010). According to Rao and Ballard (Rao and Ballard, 1999), Bayesian inference (Bayes, 1763) is a natural candidate for a generative statistical model. Within this context, Bayesian inference estimates the probability that a particular state $s$ (representing a meaning or cause) exists, given an observation $o$. This is called the posterior probability $P(s|o)$, i.e., the chance of state $s$ after, or given observation $o$. It does so by multiplying two probabilities: the likelihood $P(o|s)$ of the observation $o$ given state $s$, and the PriorPrior probability $P(s)$ of state $s$ before the observation. The result is subsequently normalized by dividing by the overall probability of the observation $P(o)$ (see Figure 1):
@@
    Signal and Noise. We call the distance between the means of the goal and progress distributions on variable $s$ the Signal to the user’s brain. It describes how far, and in which direction, the user needs to travel along the scale of task outcomes to reach the goal distribution, on average. The stronger the signal, the more stimulus the task provides. However, this also means the more work has yet to be performed by the brain. The reciprocal of this signal represents the AccuracyAccuracy of the task. The standard deviation of the goal distribution is the tolerance to noise and describes the variation allowed in the goal distribution of the task. We therefore call this the Noise in the task, because it describes how much noise is allowed in the movement towards the goal for distributions to still match. The larger the noise tolerance, the less precise the user needs to be in executing the task. As such, the reciprocal of this noise represents the PrecisionPrecision of the task. If we return to the example of organizing your socks, the larger the average distance of socks, the Signal, to the drawer, the harder it is to complete the task. The larger the drawer, or Noise tolerance, the easier it will be to complete the task. Fitts’ Law naturally follows from this, but note we are working in bits rather than time. From this also follows that the relative entropy or difficulty of a task is a function of the ratio of the mean distance S to and standard deviation N of the goal. Note that this SNR is different from the one introduced by Shannon as it describes a relative signal to a goal:
@@
Here, MTMT equals movement time, and $a$ and $b$ are empirically determined constants. The equation is modeled after the Shannon-Hartley theorem in Equation 2. We can instead define a KL divergence in the form of Equation 15 in which error $E$ is minimized:
@@
Where NN is the size of the noise tolerance in the cursor (or hand) movement ($4\sigma$ of the goal distribution) and $e$ the distance from current position to the nearest edge of the goal distribution (see Figure 2). If we rewrite this as a gradient descent (i.e., derivative of Equation 18), we obtain an affine update rule that halves the distance of the progress distribution at each step, with identical time intervals $t$:
@@
Here, $C$ equals the capacity of the human sensorimotor system (a loop that includes pointing device and cursor) in bits, with $a$ and $b$ empirically determined constants in bits. $S$ is the signal in this task, and is given by the distance between the center of the goals. Because Equation 19 describes the movement control function, rather than the exact end of movement, over many reciprocal tapping trials the expected end location clusters around the center–to–center distance $S$, with a slight bias towards the near edges. NN describes the noise tolerance of the cursor (or hand) in this task, and is equal to the width of the goal distributions. The main benefit of Equation 19 is that it can be used to evaluate Bayesian surprise, in bits, during the pointing task. Empirical validation of this Bayesian underpinning of Fitts’ Law remains ongoing. Schmidt et al. (Schmidt et al., 1979) did show the variance in motor impulse in a similar task to be linearly related to the variance in the resulting distribution, i.e., the prior. Their $W_{e}$ metric may correspond to the NN in Equation 20.
@@
Where $p_{i}$ is the probability of each choice in the set. Note that as the set gets smaller, the chance of a correct answer increases. As the set gets larger, it decreases by the number of elements in the set. If we examine Equation 7 it is evident that this is equal to a discrete KL divergence for the task, in which the distributions are probability mass functions instead of proper density functions. Here, the probability of the posterior is set to one because there is one item to which the prior needs to be updated. The prior probability is set to $p_{i}$. Conveniently, the binary logarithm of 1 reduces to zero making this relative entropy equal to the total entropy of the task. This means the only relevant parameter in this law is the noise in the prior probabilities: the knowledge the user has at the beginning of the task. According to Wu et al. (Wu et al., 2018), evidence from MRI scans of the brain indeed suggests that the narrowing of the selection from the prior to the observation follows step-wise Bayesian posterior updates. According to them, the cognitive control network in the brain eliminates half the options at each Bayesian update, yielding a binary search with a logarithmic response time. This means Hick-Hyman Law can be rewritten as a capacity function where noise NN is a discrete probability that is the reciprocal of the number of choices n:
@@
We designed a controlled car following task in a driving simulator to collect driving data in target avoidance conditions with varying noise-to-signal ratios. Our goal with this experiment was to model the drivers’ information load using our equations to gauge the amount of information used to maintain a distance $S$ when following a lead car with variation NN (see Figure 3). We focused our experimental analysis on deceleration events that may lead to collision.
$S$ and NN were varied between conditions. As in real driving, NN was not static within a condition: the lead car accelerated and decelerated to different speeds, which added a normally distributed noise NN to each distance condition. This allowed us to capture participants’ handling of the car in a range of driving difficulty levels defined by the constituting $N/S$ ratios. We hypothesized that driver’s information capacity during decelerations would be logarithmic with the noise-to-signal ratio, with a mean capacity $b$, in bits of information per bit of difficulty. Because driver information capacity is logarithmic, but the amount of information generated is quadratic, at capacity $b$, the amount of information will outstrip the driver’s ability to process. We hypothesized that this will lead to an error statistic that follows the KL divergence in Equation 32.
@@
We used a within-subject experiment where the task was to follow the lead car at a specified distance $S$. We then varied the speed of the lead car with variation NN. Each participant experienced 12 different task conditions with two factors $S$ and NN.
@@
To constitute each distribution, for each level of NN, the lead car randomly decelerated 20 times and randomly accelerated 20 times, although always keeping within the level’s range of speed values. Each trial within a level of NN comprised a new speed adjustment, which occurred approximately every two seconds. We randomized the presentation order of conditions. We also randomized the presentation order of acceleration or deceleration trials, with accelerations never occurring within a deceleration trial or vice versa.
We measured the instantaneous noise-to-signal ratio (change in distance by lead car $\Delta S$, i.e., some sample of NN; mean distance to lead car $\overline{S}$, per trial), and pedal activity by the driver. As will be detailed further, our dependent variables were the amount of information processed during a trial (KL divergence), and error (collisions). Together, the 40 trials created a condition simulating a continuous driving task of 2 minutes.
@@ Results
In our experiment, lead car accelerations never resulted in a collision. Given our interest in the relationship between target avoidance and the risk of collision (error), we constrained our analysis to trials where participants responded to a deceleration of the lead car by braking or by releasing the accelerator pedal. We used 3 different levels of precision for noise-to-signal measurements in our analysis. Each used a different confidence interval for defining NN.
@@
    At the trial level of precision, NN was defined by the actual relative backwards distance traveled by the lead car during each trial; and
@@
    Within each trial, NN was defined by the standard deviation $(\sigma)$ of a one-second time window of instantaneous distance measurements, as detailed below.
@@
As each trial involved a braking event that resulted in processing of information given an individually administered $\Delta S$, many similar $N/S$ ratios were created across conditions. We therefore used these more detailed empirical observations of noise-to-signal ratios for our log-linear regression analysis, pooled across all trials. Here, our definition of noise-to-signal ratio was $\Delta S$ per trial divided by the measured mean $S$ ($\overline{S}$) per trial, yielding the metric $\Delta S/\overline{S}$. We also used this definition for our regression analysis of error rates, as this yielded directly comparable results of error rates vs. capacity, in bits. This definition was also used for our Index of Difficulty (see Equation 36). Note that the only effect of the use of sampled $\Delta S$ values instead of some confidence interval NN is a scaling effect on $N/S$ ratios.
@@
Our first dependent variable was the amount of information processed by the driver during decelerations. To measure the instantaneous amount of information processed within a trial, we used a more detailed 1-second rolling time window to sample the distribution of distances $S$ to the lead car at 450 Hz. Our definition of noise NN here was the standard deviation ($\sigma$) of the distribution of $S$ obtained in this time window, while our definition of signal $S$ was the mean $S$ ($\overline{S}$) of this time window. First, the noise-to-signal ratio was sampled using this type of one second window at the beginning of each trial, ending at the moment the driver responded to the stimulus. We then calculated the amount of information in the stimulus by entering this first noise-to-signal ratio into Equation 32. A one-second window again sampled the noise-to-signal ratio at the end of each trial, prior to (i.e., ending at) a new deceleration or acceleration event. We calculated the remaining amount of information by entering this second noise-to-signal ratio into Equation 32.
@@
Table 1 shows the mean KL divergence values for consumed information against levels of factors NN and $S$. Since the residuals in our ANOVA model of the dependent variable differed significantly from a normal distribution (KS test, p¿0.05), we used a non-parametric Friedman test. The Friedman test showed highly significant effects of both factors NN ($\chi^{2}(2)=17.73,p<0.001$) and $S$ ($\chi^{2}(3)=40.84,p<0.001$) on the number of bits consumed by the driver (note that a two-way ANOVA produced the same results).
@@
Table 1. Mean KL (s.e.) for consumed information (in bits) for levels of factors NN and $S$.: A table of consumed information in bits of 4 levels S against 3 levels N.
@@
Figure 6 shows the relationship between error (red solid) and aggregated signal-to-noise measurements on a $\Delta S/\overline{S}$ scale across conditions, as well as on a logarithmic scale (IDID, bottom), compared with driver capacity (blue solid) and remaining capacity (black solid). The total number of collisions across all trials was 90. Error rates below or equal to a noise-to-signal ratio or IDID of 1 trended, on average, slightly below the expected 4% (at 0.05 bits or 3.4%, between 2 and 4 collisions across trials), while those above 1 increased quadratically, averaging 0.38 bits (23%), with a maximum of 0.82 bits (43% or 39 collisions across trials) at $ID=1.85$.
@@
Figure 7 shows participants’ questionnaire responses ranking difficulty per conditional $N/S$, sorted by IDID. These represent mean responses to the question “This driving task was difficult”, on a scale from 1 (Strongly Disagree) to 7 (Strongly Agree) (Lewis-Evans et al., 2010). A Friedman non-parametric analysis of variance showed participants rated the car following task as significantly more difficult with increasing IDID ($\chi^{2}(11)=116.4,p<0.0001$). Note that all participants found the task to be easy to interpret and execute, while none reported distraction by the floating bar graphic during the task.

### v5

@@ Abstract.
booktitle: Human Computer Interaction Handbook, Fourth Edition (4th ed.)
@@ Introduction
Shannon went on to define that the information capacity $C$ of the transmission of a message through an analog channel has a maximum that is the (binary) logarithm of the ratio between the signal strength $S$ and the amount of Gaussian noise $N$ in the channel, multiplied by some constant $b$, as expressed by the Shannon-Hartley Theorem:
@@
Where $b$ is some empirically derived bandwidth parameter. This equation describes the capacity of information sent through the channel in terms of the signal-to-noise ratio. Note that without the negative sign, the SNR represents the reciprocal of the probability of a transmission. A high signal allows for more uncertainty to be carried, as does a lower noise. This means the very nature of information is statistical, with underlying equations based on (e.g., normal) distributions. This is important because humans, unlike traditional computer algorithms, operate in a probabilistic environment. That is, their exact behaviour cannot be computed, however, it can be modeled using (normal) probability distributions. Using this framework, Card, Moran and Newell (Card et al., 1983) developed the Model Human Processor (MHP), a comprehensive set of heuristics for predicting human performance in simple computer tasks. Statistical elements to these models were derived from empirical observations such as those performed by Hick (Hick, 1952) and Fitts (Fitts, 1954). They demonstrated that human performance in certain tasks indeed seems to follow a similar pattern to the Shannon-Hartley Theorem: a binary logarithm of the signal-to-noise ratio in the task. E.g., Hick’s Law describes the response time $RT$ for selecting from a number of alternatives $n$, plus 1:
@@
Here, $b$ is some empirically derived constant that models the time per unit of difficulty, the log component of the task. Response time is derived by multiplying the amount of information in the task by this empirical constant. Fitts (Fitts, 1954) derived a very similar law describing the movement time $MT$ of a hand towards a target along a single dimensional trajectory $A$ with a target width of $W$ (specific equation from (MacKenzie, 1992)):
@@
Instead of framing human processing as a sequential computer processor, McClelland and Rumelhart’s parallel processing model (McClelland and Rumelhart, 1981) described perception as an interaction of top-down (cognitive/semantic) and bottom-up (sensory) elements. Rao and Ballard (Rao and Ballard, 1999) further elucidated this approach by suggesting that the brain builds up statistical knowledge of its environment, i.e., incoming messages, over time, and that this model is used generatively, i.e., top-down, to predict the subsequent messages. The error between the prediction and the messages would lead to an update of the prediction. Prediction is convenient for more than one reason: not only does it allow humans to plan actions in their environment based on relevant information, it serves to compress the information content of their environment by processing only those messages that are surprising: i.e., not, or insufficiently predicted. Predictive coding theory explains how the human brain can work so efficiently and explains the mechanisms behind attention and the scheduling of cognitive resources based on the amount of surprise in their predictions (Baldi and Itti, 2010). According to Rao and Ballard (Rao and Ballard, 1999), Bayesian inference (Bayes, 1763) is a natural candidate for a generative statistical model. Within this context, Bayesian inference estimates the probability that a particular state $s$ (representing a meaning or cause) exists, given an observation $o$. This is called the posterior probability $P(s|o)$, i.e., the chance of state $s$ after, or given observation $o$. It does so by multiplying two probabilities: the likelihood $P(o|s)$ of the observation $o$ given state $s$, and the $Prior$ probability $P(s)$ of state $s$ before the observation. The result is subsequently normalized by dividing by the overall probability of the observation $P(o)$ (see Figure 1):
@@
    Signal and Noise. We call the distance between the means of the goal and progress distributions on variable $s$ the Signal to the user’s brain. It describes how far, and in which direction, the user needs to travel along the scale of task outcomes to reach the goal distribution, on average. The stronger the signal, the more stimulus the task provides. However, this also means the more work has yet to be performed by the brain. The reciprocal of this signal represents the $Accuracy$ of the task. The standard deviation of the goal distribution is the tolerance to noise and describes the variation allowed in the goal distribution of the task. We therefore call this the Noise in the task, because it describes how much noise is allowed in the movement towards the goal for distributions to still match. The larger the noise tolerance, the less precise the user needs to be in executing the task. As such, the reciprocal of this noise represents the $Precision$ of the task. If we return to the example of organizing your socks, the larger the average distance of socks, the Signal, to the drawer, the harder it is to complete the task. The larger the drawer, or Noise tolerance, the easier it will be to complete the task. Fitts’ Law naturally follows from this, but note we are working in bits rather than time. From this also follows that the relative entropy or difficulty of a task is a function of the ratio of the mean distance S to and standard deviation N of the goal. Note that this SNR is different from the one introduced by Shannon as it describes a relative signal to a goal:
@@
Here, $MT$ equals movement time, and $a$ and $b$ are empirically determined constants. The equation is modeled after the Shannon-Hartley theorem in Equation 2. We can instead define a KL divergence in the form of Equation 15 in which error $E$ is minimized:
@@
Where $N$ is the size of the noise tolerance in the cursor (or hand) movement ($4\sigma$ of the goal distribution) and $e$ the distance from current position to the nearest edge of the goal distribution (see Figure 2). If we rewrite this as a gradient descent (i.e., derivative of Equation 18), we obtain an affine update rule that halves the distance of the progress distribution at each step, with identical time intervals $t$:
@@
Here, $C$ equals the capacity of the human sensorimotor system (a loop that includes pointing device and cursor) in bits, with $a$ and $b$ empirically determined constants in bits. $S$ is the signal in this task, and is given by the distance between the center of the goals. Because Equation 19 describes the movement control function, rather than the exact end of movement, over many reciprocal tapping trials the expected end location clusters around the center–to–center distance $S$, with a slight bias towards the near edges. $N$ describes the noise tolerance of the cursor (or hand) in this task, and is equal to the width of the goal distributions. The main benefit of Equation 19 is that it can be used to evaluate Bayesian surprise, in bits, during the pointing task. Empirical validation of this Bayesian underpinning of Fitts’ Law remains ongoing. Schmidt et al. (Schmidt et al., 1979) did show the variance in motor impulse in a similar task to be linearly related to the variance in the resulting distribution, i.e., the prior. Their $W_{e}$ metric may correspond to the $N$ in Equation 20.
@@
Where $p_{i}$ is the probability of each choice in the set. Note that as the set gets smaller, the chance of a correct answer increases. As the set gets larger, it decreases by the number of elements in the set. If we examine Equation 7 it is evident that this is equal to a discrete KL divergence for the task, in which the distributions are probability mass functions instead of proper density functions. Here, the probability of the posterior is set to one because there is one item to which the prior needs to be updated. The prior probability is set to $p_{i}$. Conveniently, the binary logarithm of 1 reduces to zero making this relative entropy equal to the total entropy of the task. This means the only relevant parameter in this law is the noise in the prior probabilities: the knowledge the user has at the beginning of the task. According to Wu et al. (Wu et al., 2018), evidence from MRI scans of the brain indeed suggests that the narrowing of the selection from the prior to the observation follows step-wise Bayesian posterior updates. According to them, the cognitive control network in the brain eliminates half the options at each Bayesian update, yielding a binary search with a logarithmic response time. This means Hick-Hyman Law can be rewritten as a capacity function where noise $N$ is a discrete probability that is the reciprocal of the number of choices n:
@@
We designed a controlled car following task in a driving simulator to collect driving data in target avoidance conditions with varying noise-to-signal ratios. Our goal with this experiment was to model the drivers’ information load using our equations to gauge the amount of information used to maintain a distance $S$ when following a lead car with variation $N$ (see Figure 3). We focused our experimental analysis on deceleration events that may lead to collision.
$S$ and $N$ were varied between conditions. As in real driving, $N$ was not static within a condition: the lead car accelerated and decelerated to different speeds, which added a normally distributed noise $N$ to each distance condition. This allowed us to capture participants’ handling of the car in a range of driving difficulty levels defined by the constituting $N/S$ ratios. We hypothesized that driver’s information capacity during decelerations would be logarithmic with the noise-to-signal ratio, with a mean capacity $b$, in bits of information per bit of difficulty. Because driver information capacity is logarithmic, but the amount of information generated is quadratic, at capacity $b$, the amount of information will outstrip the driver’s ability to process. We hypothesized that this will lead to an error statistic that follows the KL divergence in Equation 32.
@@
We used a within-subject experiment where the task was to follow the lead car at a specified distance $S$. We then varied the speed of the lead car with variation $N$. Each participant experienced 12 different task conditions with two factors $S$ and $N$.
@@
To constitute each distribution, for each level of $N$, the lead car randomly decelerated 20 times and randomly accelerated 20 times, although always keeping within the level’s range of speed values. Each trial within a level of $N$ comprised a new speed adjustment, which occurred approximately every two seconds. We randomized the presentation order of conditions. We also randomized the presentation order of acceleration or deceleration trials, with accelerations never occurring within a deceleration trial or vice versa.
We measured the instantaneous noise-to-signal ratio (change in distance by lead car $\Delta S$, i.e., some sample of $N$; mean distance to lead car $\overline{S}$, per trial), and pedal activity by the driver. As will be detailed further, our dependent variables were the amount of information processed during a trial (KL divergence), and error (collisions). Together, the 40 trials created a condition simulating a continuous driving task of 2 minutes.
@@ Results
In our experiment, lead car accelerations never resulted in a collision. Given our interest in the relationship between target avoidance and the risk of collision (error), we constrained our analysis to trials where participants responded to a deceleration of the lead car by braking or by releasing the accelerator pedal. We used 3 different levels of precision for noise-to-signal measurements in our analysis. Each used a different confidence interval for defining $N$.
@@
    At the trial level of precision, $N$ was defined by the actual relative backwards distance traveled by the lead car during each trial; and
@@
    Within each trial, $N$ was defined by the standard deviation $(\sigma)$ of a one-second time window of instantaneous distance measurements, as detailed below.
@@
As each trial involved a braking event that resulted in processing of information given an individually administered $\Delta S$, many similar $N/S$ ratios were created across conditions. We therefore used these more detailed empirical observations of noise-to-signal ratios for our log-linear regression analysis, pooled across all trials. Here, our definition of noise-to-signal ratio was $\Delta S$ per trial divided by the measured mean $S$ ($\overline{S}$) per trial, yielding the metric $\Delta S/\overline{S}$. We also used this definition for our regression analysis of error rates, as this yielded directly comparable results of error rates vs. capacity, in bits. This definition was also used for our Index of Difficulty (see Equation 36). Note that the only effect of the use of sampled $\Delta S$ values instead of some confidence interval $N$ is a scaling effect on $N/S$ ratios.
@@
Our first dependent variable was the amount of information processed by the driver during decelerations. To measure the instantaneous amount of information processed within a trial, we used a more detailed 1-second rolling time window to sample the distribution of distances $S$ to the lead car at 450 Hz. Our definition of noise $N$ here was the standard deviation ($\sigma$) of the distribution of $S$ obtained in this time window, while our definition of signal $S$ was the mean $S$ ($\overline{S}$) of this time window. First, the noise-to-signal ratio was sampled using this type of one second window at the beginning of each trial, ending at the moment the driver responded to the stimulus. We then calculated the amount of information in the stimulus by entering this first noise-to-signal ratio into Equation 32. A one-second window again sampled the noise-to-signal ratio at the end of each trial, prior to (i.e., ending at) a new deceleration or acceleration event. We calculated the remaining amount of information by entering this second noise-to-signal ratio into Equation 32.
@@
Table 1 shows the mean KL divergence values for consumed information against levels of factors $N$ and $S$. Since the residuals in our ANOVA model of the dependent variable differed significantly from a normal distribution (KS test, p¿0.05), we used a non-parametric Friedman test. The Friedman test showed highly significant effects of both factors NN ($\chi^{2}(2)=17.73,p<0.001$) and $S$ ($\chi^{2}(3)=40.84,p<0.001$) on the number of bits consumed by the driver (note that a two-way ANOVA produced the same results).
@@
Table 1. Mean KL (s.e.) for consumed information (in bits) for levels of factors $N$ and $S$.: A table of consumed information in bits of 4 levels S against 3 levels N.
@@
Figure 6 shows the relationship between error (red solid) and aggregated signal-to-noise measurements on a $\Delta S/\overline{S}$ scale across conditions, as well as on a logarithmic scale ($ID$, bottom), compared with driver capacity (blue solid) and remaining capacity (black solid). The total number of collisions across all trials was 90. Error rates below or equal to a noise-to-signal ratio or $ID$ of 1 trended, on average, slightly below the expected 4% (at 0.05 bits or 3.4%, between 2 and 4 collisions across trials), while those above 1 increased quadratically, averaging 0.38 bits (23%), with a maximum of 0.82 bits (43% or 39 collisions across trials) at $ID=1.85$.
@@
Figure 7 shows participants’ questionnaire responses ranking difficulty per conditional $N/S$, sorted by $ID$. These represent mean responses to the question “This driving task was difficult”, on a scale from 1 (Strongly Disagree) to 7 (Strongly Agree) (Lewis-Evans et al., 2010). A Friedman non-parametric analysis of variance showed participants rated the car following task as significantly more difficult with increasing IDID ($\chi^{2}(11)=116.4,p<0.0001$). Note that all participants found the task to be easy to interpret and execute, while none reported distraction by the floating bar graphic during the task.